# SurgVidLM: Towards Multi-grained Surgical Video Understanding with Large Language Model

### v1

@@ 3.2 Results
Table 2: Ablation study on StageFoucs and Multi-grained Fusion Attention.
StageFocus Multi-grained Fusion Attention Fine-grained Video Description Fine-grained Visual Reasoning
@@
2.43 2.85 3.22 2.57 0.108 0.590 0.351 0.359
2.57 2.95 3.39 2.77 0.126 0.782 0.371 0.384
2.83 3.08 3.51 2.86 0.161 0.901 0.418 0.377
Our ablation study confirms the complementary roles of StageFocus and Multi-frequency Fusion Attention, with their combination achieving the best performance across video description and visual reasoning tasks (Table 2). Using StageFocus alone reduces the correctness of information score by 0.26 in fine-grained video description, highlighting the importance of MFA for integrating global visual context. Meanwhile, StageFocus effectively preserves detailed and contextual information. These results demonstrate that both mechanisms work together to enhance accuracy and completeness.
4 Conclusions
In this work, we introduce SVU-31K, a large-scale multi-grained surgical video understanding dataset that comprises over 31K video-instruction pairs. Based on SVU-31K, we propose SurgVidLM, the first surgical video language model specifically designed for multi-grained surgical video comprehension. Our proposed StageFocus mechanism utilizes a two-stage reasoning approach to guide the model to analyze surgical procedures holistic to detail progressively. The integration of Multi-frequency Fusion Attention further enhances the model’s performance by effectively combining low- and high-frequency visual tokens, thus preserving critical information essential for fine-grained surgical scene understanding. Experimental results demonstrate that SurgVidLM significantly outperforms SOTA Vid-LLMs across various multi-grained surgical video understanding tasks. These findings underscore the potential of SurgVidLM to advance the field of surgical video understanding. Future work will focus on expanding the dataset and exploring the application of SurgVidLM in real-time surgical environments, thereby contributing to improved clinical outcomes through enhanced video analysis capabilities.

### v2

@@ IV-B Results
Fig. 3 highlights SurgVidLM’s superior performance in procedural reasoning and anatomical understanding. When identifying the pelvic plexus, SurgVidLM correctly infers that the subsequent key anatomical landmark is Denonvilliers’ fascia, closely aligning with the surgical ground truth. In contrast, VideoLLaMA-3 and Qwen2-VL tend to produce vague or tool-oriented responses, failing to capture the precise procedural progression. Similarly, when asked about the rationale for placing the incision above the Douglas pouch, SurgVidLM not only provides the correct anatomical reasoning—ensuring complete mesorectal excision and avoiding incomplete dissection—but also articulates the clinical motivation behind this choice. These qualitative results demonstrate that, beyond surface-level descriptions, SurgVidLM is capable of producing step-wise surgical objectives (mesorectal excision) and risks (avoiding tissue injury), demonstrating the effectiveness of instruction-tuning tasks like Plan, Pre-requisites, Causal, Notice, and Suggestion.
Our ablation study confirms the complementary roles of StageFocus and Multi-frequency Fusion Attention, with their combination achieving the best performance across video description and visual reasoning tasks (Table II). When only StageFocus is applied, the model already shows notable gains in correctness and temporal coherence by progressively refining global-to-local reasoning. However, this configuration still suffers from limited integration of global context into fine-grained observations, leading to incomplete detail preservation. The joint application of StageFocus and MFA achieves the best performance across all metrics: StageFocus ensures progressive refinement of surgical procedure comprehension, while MFA effectively bridges global and local representations, together enabling SurgVidLM to deliver an accurate, detail-oriented, and contextually coherent understanding of surgical videos.
TABLE II: Ablation study on StageFoucs and Multi-grained Fusion Attention.
StageFocus
Multi-grained
Fusion Attention
Fine-grained Video Description Fine-grained Visual Reasoning
@@
1.78 1.75 1.98 1.63 15.01 8.47 16.57 33.06
2.11 2.37 3.04 2.23 16.08 9.17 17.78 36.95
2.27 2.62 3.06 2.44 16.16 9.30 17.91 36.97
TABLE III: Ablation study on sampling frequency (fps).
Sampling
Frequency
Fine-grained Video Description Fine-grained Visual Reasoning
CI DO CU TU BLEU-4 CIDEr METEOR ROUGE-L
0.5 2.14 2.06 2.52 2.02 14.52 8.52 17.79 35.71
1 2.23 2.48 2.84 2.29 16.1 9.21 17.88 35.98
2 2.27 2.62 3.06 2.44 16.16 9.30 17.91 36.97
4 2.33 2.59 2.96 2.62 16.19 9.29 17.90 36.99
We also investigate the influence of different sampling frequencies on model performance, which is illustrated in Table III. Notably, the results exhibit substantial improvements when increasing the sampling frequency from 0.5 fps to 1 fps, with further gains from 1 fps to 2 fps, but no obvious improvements from 2 fps to 4 fps. We attribute this to two primary factors. Firstly, when the sampling rate increases from a lower value, the newly sampled frames are temporally farther from their neighbors, and therefore can introduce considerable incremental information. As the sampling frequency continues to increase from a higher value, the newly sampled frames become increasingly similar to the adjacent frames, yielding limited additional information while increasing redundancy. Second, surgical videos contain transient noises, such as camera shake and motion artifacts from instruments, and the higher sampling rates increase the probability of capturing these noises, diminishing the benefit of denser sampling. Therefore, continuously increasing sampling frequency does not produce linear improvements while increasing the computational costs. To strike a balance between capturing detailed information and computational efficiency, we use 2 fps as the sampling frequency for fine-grained videos as an optimal trade-off.
V Conclusions
In this work, we introduce SVU-31K, a large-scale multi-grained surgical video understanding dataset that comprises over 31K video-instruction pairs. Based on SVU-31K, we propose SurgVidLM, the first surgical video language model specifically designed for multi-grained surgical video comprehension. Our proposed StageFocus mechanism utilizes a two-stage reasoning approach to guide the model to analyze surgical procedures from holistic to detail progressively. The integration of Multi-frequency Fusion Attention further enhances the model’s performance by effectively combining low- and high-frequency visual tokens, thus preserving critical information essential for fine-grained surgical scene understanding. Experimental results demonstrate that SurgVidLM significantly outperforms SOTA Vid-LLMs across various multi-grained surgical video understanding tasks. These findings underscore the potential of SurgVidLM to advance the field of surgical video understanding. Future work will focus on expanding the dataset and exploring the application of SurgVidLM in real-time surgical environments, thereby contributing to improved clinical outcomes through enhanced video analysis capabilities.